# Work-From-Home and Privacy: What Do Workers Face and What are They Doing About it?

### v1

@@ 3.4. Not Adopting Measures
Oblivious to measures
Many participants who said they did not take any measures appeared, based on their answers, to actually be taking preventative steps, but did not realize that those steps were a type of preventative measure we were asking about. For example, P_040 said they did not take measures to prevent data scenarios, however, when asked about the reason, she elaborated: “I only use my work computer for work related emails and web browsing so I do not feel the need for protective measures.” Similarly, in audio scenarios, P_131 noted that she “found a quiet space/designated area, so no need for protective measures.” The steps taken by these participants, using a separate computer for work purposes or using a quiet or designated space, are indeed measures to prevent privacy-invasive scenarios.
Understanding work environment
An understanding and friendly work environment was mentioned across all scenarios as a reason to not adopt measures, notably in relation to autonomy scenarios, where this theme also includes flexible or no work policies that restrict worker’s autonomy as summarized by P_168: “My organization has always been understanding that we might need to turn off our microphone or camera. There’s been no negative repercussions around doing so,” and to a less extent in other scenarios: audio, video, and data.
Nothing can be done
@@
Have nothing to hide
@@ 3.5. Characterizing Privacy-Invasive Scenarios
To gain more understanding of the factors that make a privacy-invasive scenario uncomfortable, we asked participants to select one of the scenarios (the most memorable) that made them feel uncomfortable and asked them detailed follow-up questions about it. Through these detailed questions, we identified some notable patterns of scenarios that caused discomfort as we will elaborate next (see Table 10 in Appendix A for further details). For participants who did not report any scenario that made them feel uncomfortable (they either felt comfortable or never experienced them), we asked them to describe an incident that would make them feel uncomfortable and why.
@@ 3.5.1. Reasons for Discomfort
We first look at the scenarios that the participants selected as the most memorable scenario that made them feel uncomfortable. From the 140 participants who reported one or more scenarios that made them feel uncomfortable, we find that the most selected memorable scenarios tended to be from the audio category (74/140), and the least selected scenarios are from the video category (13/140). See Table 9 in Appendix A for further details.
Table 4. Frequencies of themes that form the reasons for discomfort when experiencing privacy-invasive scenarios while working from home.
@@
Our qualitative analysis of the open-ended question about why participants felt uncomfortable (Q.21) revealed four overarching themes for the causes of discomfort: affecting the flow of the meeting, affecting the worker’s professional image, feeling a lack of control over the incident, and fear of consequences. Table 4 lists the frequencies of the identified themes that form the reasons that caused discomfort when experiencing privacy-invasive scenarios while working from home. Audio scenarios have the highest frequencies in three out of the four themes.
@@
For the 74 participants who did not experience any scenario that made them feel uncomfortable, the described scenarios that would make them feel uncomfortable were centered around the workers themselves and other adults, with the top four scenarios related to: their video 24/74 , their voice 11/74, an adult video 11/74, and an adult voice 10/74. Most of the scenarios contained exacerbated details that our scenarios did not include. For example, 26/74 participants provided scenarios related to inappropriateness, mostly related to inappropriately dressed or undressed, such as “If I inadvertently shared my camera and was not dressed appropriately for work.” (P_155) and “If i was in the nude thinking my camera was off but it actually wasn’t” (P_143), as well as inappropriate behaviors, or conversations.
@@
Looking at the reasons that would make these anticipated scenarios uncomfortable, they are the same reasons that were expressed by those who experienced our scenarios, but with different priorities. Here, we observe that reasons related to interrupting the flow of the meeting are barely mentioned, and most of the reasons are centered around affecting the worker’s professional image (28/74), feeling a lack of control over the incident (17/74), and fear of consequences (6/74).
3.5.2. When Discomfort Escalates to Harm.
@@
Most participants 104/140 said the scenario did not escalate to cause harm to self or others, while 36/140 said it caused some amount of harm (35 chose “small amount of harm” and one “large amount of harm”). Out of the 36 participants who reported harm, 24/36 participants reported audio scenarios, 6/36 data, compared to only 4/36 autonomy, and only 2/36 video scenarios.
@@
Out of the 36 who reported harm, 34 said that the harm was “Psychological harm,” one participant reported “Financial harm” in addition to psychological, one reported “Physical harm,” and one selected “Other” and specified it was a “Reputational harm” (which fits under psychological harm but we keep answers as reported). When we asked them to explain how it caused the harm, the most recurrent themes were related to: feeling embarrassed mentioned by 11 participants, unprofessional (4), fear of incident recurrence or paranoia (3), feeling ashamed (2), and anxious (2). P_187 explained the psychological harm: “It made me paranoid of the situation repeating and constantly having to make sure to an almost extreme extent that everything is in order before I use my camera.” The one financial harm was due to one month’s suspension from work, the physical harm was in relation to the participant’s loud dog barking that he could not stop and explained the harm to others as “sudden loudness … which hurt their ears” (P_105), while the reputational harm was because the incident “caused some within my organization to talk” (P_208).
@@
The harm has little relation to whether there were manager(s) presented in the meeting or not. Out of the 36 participants who reported harm, nearly half (17/36) reported that there were managers in the meeting, while the remaining did not.
The Most Shared Device and Recording.
@@
Participants reported microphone was shared more often in the incidents that caused them discomfort (115/140), compared to 82/140 who shared the camera, and 28/140 who shared the screen. In over half of the cases (88/140), participants said those incidents were not recorded, suggesting that even non-recorded incidents still cause discomfort to a large extent.
Recurrence and Location.
@@
Our results suggest that whether the incident occurred once or multiple times, and whether the incident occurred in an office room or other rooms, have little impact on discomfort. That is, nearly half of the participants (71/140) reported that the incident that caused them discomfort occurred once, while a total of 69/140 reported more than once (“A few times” or “Repeatedly”). Similarly, 59/140 reported that the incidents occurred in an office room, while a total of 81/140 in a room other than an office.
4. Summary of KEY Results and Future Work
In what follows, we summarize our key findings and raise open questions that motivate future work.
@@ 4.1. WFH Discomfort is Unlikely to Cause Harm, yet Harmed Minority Should Not be Ignored
WFH privacy invasive scenarios are prevalent. Almost all participants experienced at least one scenario. More than half of the participants felt uncomfortable in at least one of the presented scenarios. When asked about harm, most participants who experienced discomfort did not find the discomfort caused harm to themselves or others (74.3%). Almost all of those who reported that the discomfort had escalated to harm said that the harm was psychological and minor. While only 25.7% of participants who experienced discomfort (16.8% of all participants) said it escalated to harm, this is still a considerable minority that should not be ignored.
Thus, organizations that intend to completely shift their work setup to WFH as witnessed by some businesses following the COVID-19 pandemic (Christian, 2023; Lufkin, 2022), need to consider this harmed minority.
4.2. WFH Policies that Restrict the Worker’s Autonomy Should be Avoided
@@
Our results show that autonomy scenarios that limit the subject’s autonomy, such as prohibiting them from turning off their camera or microphone during a work-related call or meeting are the least prevalent. However, when they are experienced, these scenarios have the highest discomfort rate compared to other scenarios. Moreover, as shown through our participants, many workers reported measures that violated or would violate these rules when needed, despite the stated rules in the autonomy scenarios, such as turning off the camera or muting the microphone to protect their privacy, which may result in tension and conflicting situations.
We suggest that employers refrain from policies that limit the worker’s autonomy such as disallowing them to turn off their cameras or microphones. Our results show that adherence to such policies may not be always possible, and there are no practical measures to prevent emergencies that may require workers to turn off their camera or microphone.
@@
4.3. Audio Scenarios are More Related to Discomfort and Harm than Video Scenarios
@@
Our results show that audio scenarios are the most experienced scenarios among our participants. We note that the prevalence of audio scenarios might be in part due to the frequency with which participants were sharing their audio as opposed to video or screen sharing. Regardless, audio scenarios have the second highest ratio of discomfort, after the much less frequently experienced autonomy scenarios. Comparing the discomfort ratio in the audio vs. video scenarios respectively (46.2% vs. 28.8%), considerably more audio scenarios caused discomfort than video scenarios. Additionally, when we asked participants to select the most memorable uncomfortable scenario from the list of all reported uncomfortable scenarios, the four most reported scenarios were audio scenarios. On the other hand, the least four reported memorable scenarios were video scenarios. Furthermore, looking at the reasons that made the memorable scenarios uncomfortable, we find that audio scenarios appeared more frequently in relation to interrupting or distracting the meeting and affecting the worker’s professional image than video and other scenarios. Audio scenarios appeared considerably more than video scenarios in scenarios that escalated to cause harm (24/36 audio vs. 2/36 video).
@@
The aforementioned findings show that audio scenarios are more prevalent, more related to discomfort and harm, suggesting that they are more difficult to prevent than video scenarios. Efforts to develop and improve easy-to-use use smart measures to prevent audio privacy-invasive scenarios may be a worthwhile endeavor.
4.4. “Smart” Measures are Not Widely Adopted.
@@
As our results show, applying “smart” privacy-protective measures through software/hardware settings or features that would automate the process (e.g. set the microphone to mute by default instead of manually muting the microphone), or using technical means to replace manual efforts (e.g. use virtual or blurred background instead of manually clearing the physical background), are not widely adopted measures among our participants compared to manual measures. Although we equally primed participants to the two types of measures when we asked them to describe the measures they took “including any new software you started using, changes to software settings, changes to your physical workspace, or anything else,” still, only a minority of participants who took measures in each category reported smart measures: 16/122 (13.1%) participants in voice scenarios, 25/104 (24.0%) in video, 8/95 (8.4%) in computer/browsing data, while only 2/48 (4.2%) mentioned smart measures to prevent autonomy scenarios. One possible explanation for such phenomena is a lack of awareness of some smart measures that are applied by default. For example, some conference tools such as Zoom have the “Zoom background noise removal” set to “Auto (automatically adjust noise suppression)” by default. There is no indicator signaling the activation status of this setting in Zoom. Thus some participants may be oblivious to it.
@@
It remains unclear what are the reasons for the low usage of smart measures in general. Is it a lack of need, awareness (of their existence or activation status), convenience, usability, skills to use them, or lack of trust in the effectiveness of smart measures in conference tools? We leave this question to future work.
4.5. Do Smart Measure Visualization/Indicators Matter?
@@
While smart measure adoption was low in general, from this study, we observe that smart measures to prevent video scenarios such as virtual or blurred backgrounds which are reflected on the worker’s and the meeting participants’ screens were more adopted than other types of smart measures that are not reflected on the meeting’s screen such as noise cancellation features. Our results show that 25/104 (24.0%) participants in video scenarios mentioned smart measures compared to 16/122 (13.1%) participants in audio scenarios.

### v3

@@ 4.4. Not Adopting Measures
Oblivious to measures.
Many participants who said they did not take any measures appeared (based on their answers) to be taking preventative steps, but did not realize that those steps were a type of preventative measure we were asking about. For example, P_040 said they did not take measures to prevent data scenarios, however, when asked about the reason, she elaborated: “I only use my work computer for work-related emails and web browsing so I do not feel the need for protective measures.” Similarly, in audio scenarios, P_131 said that she “found a quiet space/designated area, so no need for protective measures.” The steps taken by these participants, using a separate computer for work purposes or using a quiet or designated space, are indeed measures to prevent privacy-invasive scenarios.
Understanding work environment.
An understanding and friendly work environment was mentioned across all scenarios as a reason to not adopt measures, notably in relation to autonomy scenarios, where this theme also includes flexible or no work policies that restrict worker’s autonomy as summarized by P_168: “My organization has always been understanding that we might need to turn off our microphone or camera. There’s been no negative repercussions around doing so,” and to a lesser extent in other scenarios: audio, video, and data.
Nothing can be done.
@@
Have nothing to hide.
@@ 4.5. Characterizing Privacy-Invasive Scenarios
To better understand the factors that make a privacy-invasive scenario uncomfortable, we asked participants to select one of the scenarios (the most memorable) that made them feel uncomfortable and asked them detailed follow-up questions about it. Through these detailed questions, we identified some notable patterns of scenarios that caused discomfort as we will elaborate next (see Table 12 in Appendix B for further details). For participants who did not report any scenario that made them feel uncomfortable (they either felt comfortable or had never experienced any of the presented scenarios), we asked them to describe an incident that would make them feel uncomfortable and why.
@@ 4.5.1. Reasons for Discomfort
We first look at the scenarios that the participants selected as the most memorable scenario that made them feel uncomfortable. Of the 140 participants who reported one or more scenarios that made them feel uncomfortable, we find that the most selected memorable scenarios tended to be from the audio category (74/140), and the least selected scenarios are from the video category (13/140). See Table 11 in Appendix B for further details.
Table 4. Frequencies of the identified themes that form the reasons for discomfort when experiencing privacy-invasive scenarios while working from home.
@@
Our qualitative analysis of the open-ended question about why participants felt uncomfortable (Q.21) revealed four overarching themes for the causes of discomfort: affecting the flow of the meeting, affecting the worker’s professional image, feeling a lack of control over the incident, and fear of consequences. Table 4 lists the frequencies of the identified themes that form the reasons that caused discomfort when experiencing privacy-invasive scenarios while working from home. The audio scenarios have the highest frequencies in three of the four themes.
@@
For the 74 participants who did not experience any scenario that made them feel uncomfortable, the described scenarios that would make them feel uncomfortable were centered on the workers themselves and other adults, with the top scenario categories related to [^3] Since the scenarios here were provided in response to an open-ended question and not on a list as those who said they experienced at least one scenario that made them feel uncomfortable, some participants provided scenarios that contained multiple categories. Thus, scenario categories in this section are non-exclusive and the counts represent the mentioned scenario categories and not the participants.: their video (24), their voice (11), an adult’s video (11), and an adult’s voice (10). Most of the scenarios contained exacerbated details that our scenarios did not include. For example, 26 participants provided scenarios related to inappropriateness, mainly related to being inappropriately dressed or undressed, such as “If I inadvertently shared my camera and was not dressed appropriately for work.” (P_155) and “If i was in the nude thinking my camera was off but it actually wasn’t” (P_143), as well as inappropriate behaviors or conversations.
@@
Looking at the reasons that would make these anticipated scenarios uncomfortable, they are the same reasons that were expressed by those who experienced our scenarios, but with different priorities. Here, we observe that reasons related to interrupting the flow of the meeting are barely mentioned (5), and most of the reasons are centered around affecting the worker’s professional image (28/74), feeling a lack of control over the incident (17/74), and fear of consequences (6/74).
4.5.2. When Discomfort Escalates to Harm
@@
Most of the participants 104/140 said the scenario did not escalate to cause harm to self or others, while 36/140 said it caused some amount of harm (35 chose “small amount of harm” and one “large amount of harm”). Of the 36 participants who reported harm, 24/36 participants reported audio scenarios, 6/36 data, compared to only 4/36 autonomy, and only 2/36 video scenarios.
@@
Out of the 36 who reported harm, 34 said that the harm was “Psychological harm,” one participant reported “Financial harm” in addition to psychological, one reported “Physical harm,” and one selected “Other” and specified it was a “Reputational harm” (which fits under psychological harm but we keep answers as reported). When we asked them to explain how it caused the harm, the most recurring themes were related to: feeling embarrassed mentioned by 11 participants, unprofessional (4), fear of incident recurrence or paranoia (3), feeling ashamed (2), and anxious (2). P_187 explained the psychological harm: “It made me paranoid of the situation repeating and constantly having to make sure to an almost extreme extent that everything is in order before I use my camera.” The one financial harm was due to a one-month suspension from work, the physical harm was in relation to the participant’s loud dog barking that he could not stop and explained the harm to others as “sudden loudness … which hurt their ears” (P_105), while the reputational harm was because the incident “caused some within my organization to talk” (P_208).
@@
The harm has little relation to whether there were manager(s) presented in the meeting or not. Of the 36 participants who reported harm, almost half (17/36) reported that there were managers at the meeting, while the remaining did not.
4.5.3. The most shared device and recording
@@
Participants reported microphone was shared more often in incidents that caused them discomfort (115/140), compared to 82/140 who shared the camera, and 28/140 who shared the screen. In more than half of the cases (88/140), participants said that those incidents were not recorded, suggesting that even non-recorded incidents still cause discomfort to a large extent.
4.5.4. Recurrence and location
@@
Our results suggest that whether the incident occurred once or multiple times, and whether the incident occurred in an office room or in other rooms, have little impact on discomfort. That is, nearly half of the participants (71/140) reported that the incident that caused them discomfort occurred once, while a total of 69/140 reported more than once (“A few times” or “Repeatedly”). Similarly, 59/140 reported that the incidents occurred in an office room, while a total of 81/140 in a room other than an office.
5. Discussion
This study aims to provide detailed insights of the privacy-invasive scenarios that workers face while working from home. In what follows, we summarize our key findings, provide practical recommendations, and raise open questions that can motivate future work.
@@ 5.1. WFH Discomfort is Unlikely to Cause Harm, Yet Harmed Minority Should Not be Ignored
WFH privacy-invasive scenarios are prevalent. Almost all participants experienced at least one scenario. More than half of the participants felt uncomfortable in at least one of the presented scenarios. When asked about harm, most of the participants who experienced discomfort did not find the discomfort caused harm to themselves or others (74.3%). Almost all of those who reported that the discomfort had escalated to harm said that the harm was psychological and minor. While only 25.7% of participants who experienced discomfort (16.8% of all participants) said that it escalated to harm, this is still a considerable minority that should not be ignored.
Our results suggest that the privacy-invasive scenarios that workers experience while working from home often do not cause harm. Nevertheless, organizations that intend to enforce WFH either in full or hybrid mode need to consider that there might be a harmed minority. One possible suggestion to address this is to involve workers in discussions about WFH policy, listen to the needs of those who might be harmed, and consider ways to mediate this.
5.2. Rethink WFH Policies that Restrict the Worker’s Autonomy
@@
Our results show that autonomy scenarios that limit the subject’s autonomy, such as prohibiting them from turning off their camera or microphone during a work-related call or meeting, are the least prevalent scenarios. However, when they are experienced, these scenarios have the highest discomfort rate compared to other scenarios. Moreover, as shown through our participants, many workers reported measures that violated or would violate these rules when needed, despite the stated rules in the autonomy scenarios such as turning off the camera or muting the microphone to protect their privacy, which may result in tension and conflicting situations between workers and employers.
One might argue that workers should have no expectations for privacy on employer’s owned devices and that such autonomy-invading policies are set to protect organizations (SHRM, 2018), e.g. to prevent workers from working in public spaces that can expose confidential information. However, our results show that adherence to such policies may not be always possible, and there are no practical measures to prevent emergencies that may require workers to turn off their camera or microphone. Thus, we suggest that employers rethink policies that limit the worker’s autonomy such as disallowing them to turn off their cameras or microphones. For example, we suggest employers discuss and identify the exceptional situations where workers can be exempted from these autonomy-restricting policies, which is inline with Emami-Naeini’s recommendation (Emami-Naeini et al., 2021). This may release the tension that such policies impose on worker as we observed in some participants’ comments.
@@
5.3. Audio Privacy-Invading Scenarios are More Related to Discomfort and Harm than Video Scenarios
@@
Our results show that audio scenarios are the most experienced scenarios among our participants. We note that the prevalence of audio scenarios might be in part due to the frequency with which participants shared their microphone as opposed to camera or screen sharing. Regardless, audio scenarios have the second highest discomfort ratio, after the much less frequently experienced autonomy scenarios. When comparing the discomfort ratio in all audio vs. all video scenarios, respectively 239/517 (46.2%) vs. 98/340 (28.8%), considerably more audio scenarios caused discomfort than video scenarios (see Table 9 in Appendix B for details). Additionally, when we asked participants to select the most memorable uncomfortable scenario from the list of all reported uncomfortable scenarios, the four most reported scenarios were audio scenarios. On the other hand, the least four reported memorable scenarios were video scenarios. Furthermore, looking at the reasons that made the memorable scenarios uncomfortable, we find that audio scenarios appeared more frequently in relation to interrupting or distracting the meeting and affecting the worker’s professional image than video and other scenarios. Audio scenarios appeared considerably more than video scenarios in scenarios that escalated to cause harm (24/36 audio vs. 2/36 video).
@@
The aforementioned findings show that audio scenarios are more prevalent, and more related to discomfort and harm, suggesting that they are more difficult to prevent than video scenarios. Future work can better understand the reasons that made audio privacy-invasive scenarios more prevalent, and associated to discomfort and harm compared to video scenarios.
5.4. “Smart” Measures are NOT Widely Adopted Compared to Manual Measures
@@
Our results show that applying “smart” privacy-protective measures through software/hardware settings or features that would automate the process (e.g. set the microphone to mute by default instead of manually muting the microphone) or using technical means to replace manual efforts (e.g. use virtual or blurred background instead of manually clearing the physical background) are not widely adopted measures among our participants compared to manual measures. Although we equally primed participants to the two types of measures when we asked them to describe the measures they took “including any new software you started using, changes to software settings, changes to your physical workspace, or anything else,” still, only a minority of participants who took measures in each category reported smart measures: 16/122 (13.1%) participants in voice scenarios, 25/104 (24.0%) in video, 8/95 (8.4%) in data, while only 2/48 (4.2%) mentioned smart measures to prevent autonomy scenarios.
One possible explanation for such phenomenon is a lack of awareness of some smart measures from the first place. For example, Li et al.’s study on the reasons that drive students not to share their cameras in classrooms found that many students were concerned about the appearance of their background or the people around them, suggesting that students lack awareness about smart features on conferencing tools such as virtual background (Li et al., 2022). Another reason is a lack of awareness that some features are applied by default. For example, some conference tools such as Zoom have the “Zoom background noise removal” set to “Auto (automatically adjust noise suppression)” by default. There is no indicator that signals the activation status of this setting in Zoom. Thus, some participants may be oblivious to it.
@@
It remains unclear what are the reasons for the low usage of smart measures in general. Is it a lack of need, awareness (of their existence or default activation status), convenience, usability, discoverability, skills to use them, or lack of trust in the effectiveness of smart measures in conference tools? We leave these questions to future work.
5.5. Do Smart Measures Visualization/Indicators Matter?
@@
While smart measure adoption was low in general, from this study we observe that smart measures to prevent video scenarios such as virtual or blurred backgrounds which are reflected on the worker’s and the meeting participants’ screens were more adopted than other types of smart measures that are not reflected on the meeting’s screen such as noise cancelation features. Our results show that 25/104 (24.0%) participants in video scenarios mentioned smart measures compared to 16/122 (13.1%) participants in audio scenarios.
@@
5.6. Research Questions Revisited
In what follows we summarize key findings for our main research questions.
  • How prevalent is experiencing privacy-invasive scenarios among WFH workers?
    Our results show that WFH privacy-invasive scenarios are prevalent among our participants. Almost all participants (93.9%) reported that they have experienced at least one of the privacy-invasive scenarios we presented to them.
  • To what extent does experiencing these scenarios cause discomfort?
    More than half of the participants (65.4%) felt uncomfortable during at least one of the scenarios they experienced.
  • Does discomfort escalate to cause harm to workers or others?
    Most of the participants who experienced discomfort in privacy invasive scenarios said that the discomfort did not cause harm to them or others (74.3%).
  • What type of harm do workers experience?
    Almost all the participants who reported that the discomfort they experienced escalated to cause harm to them or other said the harm was psychological (34/36) and minor (35/36).
  • What types of scenarios are associated with the highest reported discomfort?
    Our results show that scenarios that restrict worker autonomy, such as prohibiting them from turning off the camera or microphone are associated with the highest reported discomfort when experienced (75%). Participants often violated or would violate the employer’s autonomy-restricting rules to protect their privacy.
  • What protective measures do some workers take?
    Most participants reported manual measures such as manually covering the camera and turning on/off the microphone. Smart measures (software/hardware settings or features to prevent privacy-invasive scenarios, e.g. virtual background and noise cancelation features) are not widely adopted compared to manual measures. It should be noted that when smart measures are used, more participants reported using settings or features to prevent video scenarios (24.0%) than audio (13.11%), data (8.4%), and autonomy (4.2%) scenarios.
  • Why do some workers choose not to take protective measures?
    Participants reported a variety of reasons for not taking protective measures across all scenario categories with varying rates. The reported reasons include: never experiencing the scenario, having an understanding work environment, feeling that there is nothing they can do to prevent the scenario, or not feeling bothered about experiencing it. Interestingly, some participants were oblivious to measures they were already taking that may help prevent privacy-invasive scenarios. That is, they reported not taking measures, but their open-ended answers to the reasons for not doing so revealed that they are taking measures such as using a separate device for work and using a quiet or designated space for working from home.
5.7. Limitations
First, we used the Prolific (Prolific, 2024) crowdworking platform to recruit participants, and Prolific workers are known to be more technically skilled than the average population. However, Prolific has been shown to provide reasonably generalizable results in security and privacy research (Tang et al., 2022). Moreover, we screened for Prolific users who were not performing their reported jobs through crowdsourcing platforms such as Prolific and MTurk. This may have skewed our participants towards a certain type of Prolific users. Second, we used self-reported surveys, which are naturally prone to recall bias and social desirability effects. To mitigate recall bias, we asked participants to answer based on their experiences within the last three years. Moreover, in the open-ended questions when we asked participants what protective measures they took, we equally nudged participants to mention manual and smart measures. However, participants may still have trouble accurately recalling incidents or measures within a three-year time frame. Third, we acknowledge a limitation in the follow-up detailed questions on the scenarios that caused discomfort, where we only followed up with the participants about one scenario they reported as the most memorable scenario that made them feel uncomfortable. However, asking participants follow-up questions about multiple scenarios would likely result in survey fatigue that would affect the quality of the data. Fourth, we did not ask the participants what conference software they use. Different software may have different default settings and features. Moreover, some participants may not be fully aware of their software’s default settings and features. Fifth, our study does not consider potential nuanced differences between different types of jobs. This can have implications for the type of rules imposed by employers and the consequences of privacy invasions for workers. Finally, our sample is limited to participants from the US. Privacy laws and social norms can vary between jurisdictions, which needs to be considered when interpreting our results outside the US context. Even within the US, privacy regulations and labor laws vary somewhat by state, which may impact workers’ experiences.